OpenStudy (pottersheep):

find the sum for n = 7, a = 2, and r = 1/2. I get 127/128...but the answers say 127/32.

OpenStudy (pottersheep):

I use the formula Sn = (1 - 2^n) / 1-r

OpenStudy (pottersheep):

sorry a(1 - 2^n) / 1-r

OpenStudy (anonymous):

you are right

OpenStudy (pottersheep):

Really ? Thanks

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

\[r=\frac{1}{2}\]?

OpenStudy (pottersheep):

MY BAD ! a(1 - ****r^n***) / 1-r

OpenStudy (pottersheep):

n = 7, a = 2, and r = 1/2.

OpenStudy (pottersheep):

r is common ratio

OpenStudy (anonymous):

i am assuming this is \[a=2,r=\frac{1}{2}, n = 7\] in other words you have \[2+1+\frac{1}{2}+\frac{4}{4}+\frac{1}{8}+\frac{1}{16}+\frac{1}{32}\]

OpenStudy (pottersheep):

yep

OpenStudy (anonymous):

typos above, should be \[\frac{1}{4}\]

OpenStudy (pottersheep):

Yeah O got that :)

OpenStudy (pottersheep):

*I

OpenStudy (pottersheep):

so I did s7 = 2(1 - (1/2)^7)/1 - 1/2

OpenStudy (anonymous):

shouldbe \[S_n=\frac{a-ar^{n-1}}{1-r}\]

OpenStudy (anonymous):

off by a power

OpenStudy (anonymous):

damn another typo, should be \[S_n=\frac{a-ar^{n+1}}{1-r}\]

OpenStudy (pottersheep):

I my formula has a factored out at the top, same thingg I think so

OpenStudy (anonymous):

so you have \[\frac{2-(\frac{1}{2})^8}{1-\frac{1}{2}}\]

OpenStudy (pottersheep):

oh okay

OpenStudy (anonymous):

btw in any case the first to numbers you are adding are 2 and 1, so there is no way you can get \[\frac{127}{128}\] right?

OpenStudy (pottersheep):

d'awww

OpenStudy (anonymous):

sum has to be greater than 3 for sure

OpenStudy (pottersheep):

yup true

OpenStudy (anonymous):

point is not to call you out, but to remind you not to get married to a formula and forget about common sense

OpenStudy (pottersheep):

You're right, haha...thanks

OpenStudy (pottersheep):

OMG! I know what I did wrong now! I divided a fraction by a fraction wrong (always do that ugh!)....

OpenStudy (pottersheep):

Thanks for your help though ~~~~ Otherwise I would have assumed the paper was wrong :P